OpenStudy (anonymous):

d/dx [ln(5x^{3} +4)]

OpenStudy (amistre64):

what the chain rule pull out; say if this was just ln(u) ??

OpenStudy (anonymous):

okay so it would be...\[\frac{ 5x ^{2} }{ x ^{3} +4}\]

OpenStudy (amistre64):

close .. but not quite :)

OpenStudy (anonymous):

It should be 3x^{2} at the top

OpenStudy (amistre64):

recall\[\frac d{dx}(lnu)=\frac {u'}{u}\] let u = 5x^3 + 4 u' = 15x^2

OpenStudy (amistre64):

you had the right idea tho, and thats good :)

OpenStudy (anonymous):

okay, got it! Thanks :)